OpenStudy (anonymous):

whats the middle number of 60 and 68... 64 i am not sure

Parth (parthkohli):

Yes. Correct.

Parth (parthkohli):

(60 + 68)/2 = 128/2 = 64

OpenStudy (anonymous):

hm ok cus its not solving my problem with 64

OpenStudy (anonymous):

use averaging..

OpenStudy (anonymous):

ok